A diver springs upward from a board that is 4.50 m above the water. At the instant she contacts the water her speed is 14.2 m/s and her body makes an angle of 62.2 ° with respect to the horizontal surface of the water. Determine her initial velocity, both (a) magnitude and (b) direction.
Answer #1

final horizontal and vertical components of velocities are

Vx=14.2Cos62.2

Vy=-14.2Sin62.2

from kinematic equation

Vy2=Voy2-2g(y-yo)

(-14.2Sin62.2)2 = Voy2-2*9.8*(0-4.5)

Voy=8.34 m/s

Vox=Vx=14.2Cos62.2=6.623 m/s

A)

Magnitude

Vo=sqrt[Vox2+Voy2]=sqrt(6.6232+8.342)

Vo=10.65 m/s

b)

Direction

o=tan-1(Voy/Vox) =tan-1(8.34/6.623)

o=51.55o above the horizontal